Larkspur Ltd. purchased a building on January 1, 2018 for $15,360,000. Larkspur accounted for this asset using the revaluatioHere are the accounts available:

  • Accounts Payable
  • Accounts Receivable
  • Accumulated Depreciation - Buildings
  • Accumulated Depreciation - Equipment
  • Accumulated Depreciation - Leasehold Improvements
  • Accumulated Depreciation - Machinery
  • Accumulated Depreciation - Vehicles
  • Advertising Expense
  • Asset Retirement Obligation
  • Buildings
  • Cash
  • Common Shares
  • Contributed Surplus
  • Contributed Surplus - Donated Capital
  • Cost of Goods Sold
  • Deferred Revenue - Government Grants
  • Depreciation Expense
  • Donation Revenue
  • Equipment
  • Finance Expense
  • Finance Revenue
  • Gain on Disposal of Building
  • Gain on Disposal of Equipment
  • Gain on Disposal of Machinery
  • Gain on Disposal of Truck
  • Gain or Loss in Value of Investment Property
  • GST Payable
  • GST Receivable
  • Interest Expense
  • Interest Income
  • Interest Payable
  • Interest Receivable
  • Inventory
  • Investment Property
  • Land
  • Land Improvements
  • Legal Expense
  • Loss on Disposal of Building
  • Loss on Disposal of Equipment
  • Loss on Disposal of Land
  • Loss on Disposal of Machinery
  • Loss on Disposal of Truck
  • Machinery
  • Mineral Resources
  • Mortgage Payable
  • No Entry
  • Notes Payable
  • Notes Receivable
  • Office Expense
  • Owner's Drawings
  • Prepaid Expenses
  • Prepaid Insurance
  • Profit on Construction
  • Purchase Discounts
  • Purchase Returns and Allowances
  • Rent Expense
  • Repairs and Maintenance Expense
  • Revaluation Gain or Loss
  • Revaluation Surplus (OCI)
  • Revenue - Government Grants
  • Salaries and Wages Expense
  • Salaries and Wages Payable
  • Sales Revenue
  • Service Revenue
  • Supplies
  • Supplies Expense
  • Tenant Deposits Liability
  • Vehicles
Answer #1
Date Account Debit Credit
Dec 31st 2019 Buildings (Note 2) $196,000
Revaluation surplus $196,000
(being upward revaluation of building account was adjusted in the books)
Dec 31st 2019 Depreciation expense (Note 3) $519,000
Buildings $519,000
(being depreciation for the year on revalued building adjusted)
Dec 31st 2021 Revaluation surplus (Note 5) $196,000
Impairment loss $88,400
Buildings $284,400
(being impairment loss on building writtenoff to extent revaluation surplus available and balance to impairment loss account)
Dec 31st 2021 Depreciation expense (Note6) $508,061
Building $508,061
(being depreciation for the year on adjusted buildings amount)

Notes:

1. Buildings was initially recognised at $15,360,000 and life is 30 years so depreciation per year is = 15360000/30 = $1024000 per year

2.Written down value of buildings on 31st december 2019 is 14,336,000 (15360000- 2*1024000)

Revaluation surplus is 14532000-14336000 = 196000

3.Depreciation from year 3 = 14532000/ 28 (balance life) = 1038000

4.Written down value of buildings on 31st december 2021 is 13,494,000 (14532000- 2*1038000)

5.New revalued figure is 13209600 therefore impairment loss = 13494000-13209600 = 284400

this loss should be first adjusted to revaluation reserve available and balance to be charged to impairment loss

6. depreciation from year 5 = 13209600/26 = 508061